OpenStudy (anonymous):

Lim Arctan ( racine 3-x) quand x tend vers -infini

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty}\arctan(3-x)=\frac{ \pi }{ 2 }\]